# Comparing Minimal and Non-Minimal Quintessence Models to 2025 DESI Data

### v1

@@ I Introduction
where $a$ is the scale factor and the parameters $w_{0}$ and $w_{a}$ denote the equation of state and its derivative today (a=1). Note that the cosmological constant is obtained for $w_{0}=-1$ and $w_{a}=0$. While the CPL parametrization is by no means fundamental or generic, it might be a good approximation in the range of redshifts [^5] Redshift is related to the scale factor via z=1+1/az=-1+1/a. to which DESI is sensitive ($0.295\leq z\leq 2.33$).
There is a wide variety of models that can describe dynamical dark energy. The simplest ones are canonical quintessence models, where the role of dark energy is played by a scalar field $\varphi$ (with standard kinetic structure and minimal coupling to gravity) that evolves under the influence of some potential $V(\varphi)$. However, one can anticipate that this class of theories may not be the best candidates to provide an excellent fit to the DESI data. According to DESI’s results, the central values of the CPL parameters are $w_{0,\rm central}\approx-0.7$ and $w_{a,\rm central}\approx-1$, and so $w_{\text{DE, central}}(a)\approx-1.7+a$. Therefore, if we were to extrapolate the CPL parametrization all the way to the past, we would find $w_{\rm DE}<-1$ for $a<0.7$, or in other words, for $z>0.43$. This means that the Null Energy Condition (NEC) would be violated for such redshifts, but this cannot happen for minimally coupled scalar fields with canonical kinetic terms. [^6] The equation of state parameter can be written as $w_{\rm DE}=-1+(p_{\varphi}+\rho_{\varphi})/\rho_{\varphi}$. The NEC states that $p_{\varphi}+\rho_{\varphi}>0$, and thus $w_{\rm DE}>-1$. For a minimally coupled scalar field, this is always true as $p_{\varphi}+\rho_{\varphi}=\dot{\varphi}^{2}$. However, we can relax this for non-minimally coupled scalars, as we discuss later. Since $z=0.43$ lies inside DESI’s range of redshifts, one concludes that an accurate fit to the DESI data is not possible, although there could still be a modest level of compatibility. Values of $w_{0}$ and $w_{a}$ closer to the central ones may be obtained in theories where NEC violation is not forbidden. These include models of scalar fields non-minimally coupled to gravity [8, 9, 10], braneworld dark energy [11], dark matter-dark energy interactions [12, 13] (see [14] for a detailed review), fields with non-standard kinetic terms [15, 16] or modified gravity theories [17, 18].
@@ IV.2 Gravitational Coupling Constraints
Secondly, as one can easily check from the action (21), the strength of gravity is controlled by an effective gravitational constant which depends on time:
@@ V Discussion
For non-minimally scalar fields, one can arrange for an improved fit to the data. However, this implies a new fifth force and a dynamical value for the effective Newton’s gravitational coupling $G$. Both are tightly constrained by tests of gravity. For some carefully selected values, we can escape bounds from the Solar System by having the fifth force “accidentally” be suppressed today. The bound on the evolution of Newton’s coupling is able to be obeyed today, though there are moments in the history of the universe where our existing example is outside the bounds; see right panel in Fig. 8. Whether this is in tension with other bounds is less clear. From Fig. 9 this implies a shifted value of $G$ in the very early universe, including Big Bang Nucleosynthesis. However, in this example, the shift is only at the level $\sim 0.3\%$ (we have not directly included radiation in our analysis to be certain, but we anticipate the field does not evolve too much in the very early universe). It would be interesting to provide further constraints on these kinds of model in the future.

### v2

@@ I Introduction
where $a$ is the scale factor and the parameters $w_{0}$ and $w_{a}$ denote the equation of state and its derivative today (a=1). Note that the cosmological constant is obtained for $w_{0}=-1$ and $w_{a}=0$. While the CPL parametrization is by no means fundamental or generic, it might be a good approximation in the range of redshifts [^5] Redshift is related to the scale factor via z=1+1/az=-1+1/a. to which DESI is sensitive ($0.295\leq z\leq 2.33$). Other alternatives and extensions of the CPL parametrization have also been discussed in the literature (see, for instance, [8, 9, 10, 11, 12, 13, 14, 15]).
There is a wide variety of models that can describe dynamical dark energy. The simplest ones are canonical quintessence models, where the role of dark energy is played by a scalar field $\varphi$ (with standard kinetic structure and minimal coupling to gravity) that evolves under the influence of some potential $V(\varphi)$. However, one can anticipate that this class of theories may not be the best candidates to provide an excellent fit to the DESI data. According to DESI’s results, the central values of the CPL parameters are $w_{0,\rm central}\approx-0.7$ and $w_{a,\rm central}\approx-1$, and so $w_{\text{DE,central}}(a)\approx-1.7+a$. Therefore, if we were to extrapolate the CPL parametrization all the way to the past, we would find $w_{\rm DE}<-1$ for $a<0.7$, or in other words, for $z>0.43$. This means that the Null Energy Condition (NEC) would be violated for such redshifts, but this cannot happen for minimally coupled scalar fields with canonical kinetic terms. [^6] The equation of state parameter can be written as $w_{\rm DE}=-1+(p_{\varphi}+\rho_{\varphi})/\rho_{\varphi}$. The NEC states that $p_{\varphi}+\rho_{\varphi}>0$, and thus $w_{\rm DE}>-1$. For a minimally coupled scalar field, this is always true as $p_{\varphi}+\rho_{\varphi}=\dot{\varphi}^{2}$. However, we can relax this for non-minimally coupled scalars, as we discuss later. Since $z=0.43$ lies inside DESI’s range of redshifts, one concludes that an accurate fit to the DESI data is not possible, although there could still be a modest level of compatibility. [^7] In models that consider an extended parameter space, this level of compatibility can be higher (see, for instance, [16, 17]). Values of $w_{0}$ and $w_{a}$ closer to the central ones may be obtained in theories where NEC violation is not forbidden (or is just apparent). These include models of scalar fields non-minimally coupled to gravity [18, 19, 20, 21] (see also [22, 23]), braneworld dark energy [24], dark matter-dark energy interactions [25, 26, 27, 28, 29, 30] (see [31] for a detailed review), fields with non-standard kinetic terms [32, 33] or modified gravity theories [34, 35, 36, 37, 38, 39]. Other possibilities are based on non-standard equations of state for dark matter [40, 41, 42] and modified cosmologies in the framework of holographic dark energy [43] or the generalized uncertainty principle [44]. To date, the strongest evidence for a phantom crossing in the equation of state of dark energy has been reported in [45] with an approximate significance level of $5\sigma$.
@@ IV.2 Gravitational Coupling Constraints
Secondly, as one can easily check from the action (21), the strength of gravity is controlled by an effective gravitational constant which depends on time [^9] In [47, 48, 49, 50], similar models with early time variation of the gravitational constant have been analyzed and tested against observations with a focus on their potential to alleviate the Hubble tension.:
@@ IV.4 Results
Finally, we note that in our construction, nothing prevents the scalar field from dominating the energy density of the universe at redshifts higher than the one corresponding to our initial condition. The shape of the potential and the coupling to the Ricci scalar chosen in this example need to be modified to some extent in order for dark energy to be subdominant at early times, including recombination, matter-radiation equality and earlier.
@@ V Discussion
For non-minimally scalar fields, one can arrange for an improved fit to the data. However, this implies a new fifth force and a dynamical value for the effective Newton’s gravitational coupling $G$. Both are tightly constrained by tests of gravity. For some carefully selected values, we can escape bounds from the Solar System by having the fifth force “accidentally” be suppressed today. The bound on the evolution of Newton’s coupling is able to be obeyed today, though there are moments in the history of the universe where our existing example is outside the bounds; see right panel in Fig. 8. Whether this is in tension with other bounds is less clear. From Fig. 9 this implies a shifted value of $G$ in the very early universe, including Big Bang Nucleosynthesis. However, in this example, the shift is only at the level $\sim 0.3\%$ (we have not directly included radiation in our analysis to be certain, and as mentioned earlier, some modification of the theory is required to properly handle the very early universe). It would be interesting to provide further constraints on these kinds of model in the future.